# OpenPBR: Novel Features and Implementation Details

### v1

@@ 5.3 Roughening
We intend to provide better approximations in the future. For example, a derivation of the roughening due to transmission through a rough GGX dielectric interface was provided by [Belcour2018], which could provide a more physically accurate result.
@@ 5.4 Total internal reflection
We suggest as a reasonable practical approximation (following [Kutz2021]) to simply invert the IOR ratio at the coat-base boundary, i.e., replace $\eta_{\mathrm{bc}}$ as follows:
@@
Both of these schemes produce a plausible appearance that eliminates the spurious TIR. A more physically correct approach would require something similar to the Weidlich–Wilkie layering model [Weidlich2007], or the layered material model described in [Pharr2023], where the scattering through the layers is explicitly modeled via Monte Carlo simulation.
@@
Previous approaches include the microfacet-based sheen of [Estevez2017] as used by the Autodesk Standard Surface model [Georgiev2019], and the single-scattering microflake model of Adobe Standard Material [Kutz2021]. In OpenPBR, we adopt the multi-scattering volumetric microflake model of [Zeltner2022], which uses the SGGX microflake representation of [Heitz2015] and provides efficient evaluation and importance sampling via linearly transformed cosines (LTCs) [Heitz2016b].
@@
  • Volumetric foundation: A homogeneous, purely scattering layer with a fiber-like SGGX microflake phase function [Heitz2015], providing a physically plausible appearance.
@@
  • Layering model: The layer has unit optical thickness, is index-matched to the exterior medium (eliminating Fresnel at the interface), and conserves energy – all unreflected light transmits uncolored, as there is no absorption, which allows for principled layering (unlike the [Estevez2017] model).
@@
where $\mathbf{F}$ = fuzz_color, $E_{\mathrm{fuzz}}(\mu_{o},\alpha)$ (termed $R$ by [Zeltner2022]) is the reflectance at angle cosine $\mu_{o}$ given roughness $\alpha$ = fuzz_roughness $\in[0,1]$, and $D(\mu_{i}|\mu_{o},\alpha)$ is a lobe defined by linear transformations of a cosine lobe (LTCs), where the transformation matrices (and $E_{\mathrm{fuzz}}$) are fitted (over $\mu_{o}$ and $\alpha$) to a simulation of the scattering in the volumetric fuzz microflake layer. Since the LTC lobe $D$ is a normalized PDF over the hemisphere, the resulting albedo of $f_{\mathrm{fuzz}}$ is $\mathbf{F}\,E_{\mathrm{fuzz}}(\mu_{o},\alpha)$.
@@
In the model of [Zeltner2022], the optical depth is at most 1, which means the opacity at normal incidence is at most $1-\exp(-1)=0.63$, i.e., at least 36% of the light gets through. In really thick dust, the opacity could be arbitrarily close to 1. If the model were to set the optical depth to a higher value (e.g., 3), thicker fuzz could be represented, though that would make the fuzz opacity more uniform across all viewing angles. Ideally, the optical depth would be exposed as a parameter. In other words, it would be helpful to be able to control the density and thickness in order to make a fully opaque layer of dust.
@@ 7.1 Thin-wall dielectric
In thin-walled mode, the translucent-base slab can be considered an infinitesimally thin sheet of dielectric (with an embedded absorbing but non-scattering medium), with the BSDF $f_{\mathrm{dielectric}}$ on both sides. A ladder of inter-reflections occurs inside this slab, producing a reflected lobe and undeflected refracted lobe. The transmission_color can be assumed to give the transmittance through the thin sheet at normal incidence due to absorption (with transmission_depth ignored). In the smooth case, the BRDF and BTDF of this sheet can be solved exactly by summing over a geometrical series of terms containing Fresnel and absorption factors, and this can be extended to a good approximation of the rough case by appropriately roughening the transmission lobe (as described by [Kulla2017]). This model of thin-walled glass is a cheaper, much more convenient way to render windows than a finite thickness non-thin-walled mesh.
This model is similar to the ThinDielectricBxDF of PBRT, except with absorption and roughness accounted for [Kulla2017, Pharr2023]. The roughening effect of the thin wall can be modeled by considering the broadening of the incident and transmitted rays. [Belcour2018] provides a useful heuristic for this, but a more complete model would be a welcome advancement. Figure 39 shows the effect of varying the specular_roughness parameter on the appearance of a thin-wall glass material.
@@ 7.2 Thin-wall subsurface
It should also be noted that there currently isn’t a completely satisfactory model for the case of the described absorbing only thin-wall dielectric with rough boundaries, for example the ThinDielectricBxDF of PBRT [Pharr2023] simplified to the case of smooth boundaries. There are suggested heuristics [Kulla2017, Belcour2018] but a fully worked out model would be beneficial.
@@
  • There is interest in adding explicit support for a model for rendering glints [Jakob2014b, Yan2014] or flakes [Atanasov2016], which are important effects for certain uses cases, such as car paint.
  • Development (currently in collaboration with [Palmqvist2025]) of a set of example materials suitable as presets, and for demonstration of the model’s capabilities.
@@ Appendix B A mixture model example
In various practical models such as Disney’s “Principled” shader [Burley2012], Autodesk 3ds Max’s Physical Material [Andersson2016] and Autodesk Standard Surface [Georgiev2019], the model is defined as a mixture (i.e., linear combination) of BSDF/BSSRDFs corresponding to the constituent BSDF lobes of the reflection and transmission. The light transport between the layers and overall energy balance is approximated via the mix weights. Given such a representation, the integration into a renderer is relatively straightforward as the individual lobes can be importance sampled according to their albedos and combined with direct lighting estimates via standard techniques such as multiple importance sampling (MIS) [Veach1998].
@@ Appendix C Similarity theory of subsurface anisotropy
As noted, there is the following “similarity relation” [Zhao2014, Hyperion] between the scattering coefficients in media with different anisotropies $g$ and $g^{\star}$:
@@ Appendix D F82-tint model fits for a variety of real metals
We present here fits to the F82-tint model used for the metallic slab. These were computed using publicly available tabulated IOR data [Polyanskiy2023]. Example programs for computing these fits can be found here:
@@ Appendix E Implementing thin‐film iridescence from first principles
Rather than the approach of [Belcour2017] in which the thin-film interference effects were pre-integrated over color matching functions, we favour a more direct and general “locally spectral” approach that computes reflectance per light path by evaluating the full Fresnel and Airy interference stack – including complex amplitudes, polarizations, and phase shifts – at specific wavelengths sampled per path. This can begin with fixed red, green, and blue wavelengths, but better results are achieved by stochastically sampling wavelengths from approximate camera sensitivity curves (as described in Section 3.1.3). This enables convergence to neutral gray for very thick films and avoids the high-frequency color banding that fixed RGB wavelengths can produce.
The same wavelengths can also be reused to model dispersion, while all other BSDF components are free to ignore them and operate in RGB as usual. This approach uses only the Airy summation (Equation 3 from [Belcour2017]) – a method whose origins trace back to 19th-century studies of thin-plate interference by G.B. Airy and others – but requires additional per-wavelength computations and assembling the necessary formulas from multiple sources rather than a single reference. We attempt to collect the necessary equations and implementation instructions here for convenience.
The geometry of the thin film is assumed to follow the model described in [Belcour2017], as illustrated in Figure 41. Table 3 summarizes the relevant mathematical quantities. The film is taken to be applied at each microfacet of the base substrate, so can be assumed to be a flat slab with normal given by the local micronormal $\mathbf{m}$, and thickness $d$. The film is assumed to have real IOR $n_{2}$ and thickness $d$, and to be surrounded by ambient medium above with (real) IOR $n_{1}$, and substrate material with IOR $n_{3}$ (conducting or dielectric, where $n_{3}$ is complex in the conducting case). The incident light from direction $\boldsymbol{\omega}_{i}$ is at angle $\theta_{1}$ to $\mathbf{m}$, while the light refracted into the film makes angle $\theta_{2}$ with $\mathbf{m}$. Given this setup, we need to determine the albedo of the resulting configuration for incident light with wavelength $\lambda$, accounting for interference effects.
@@
Figure 41: Schematic of the thin-film model (following [Belcour2017]).
@@ 3. Reciprocity for reverse direction
To avoid re-running the Fresnel equations, one can use the following symmetry relations between the reflection and transmission amplitudes [BornWolf]:
@@ Appendix F The physics of coat darkening
We present here some formulas for the physical darkening effect of light transport within a dielectric coat. In the case of a smooth dielectric coat on top of a Lambertian substrate (the so-called “interfaced Lambertian” model), there is an exact solution available, which we describe in detail below. This was derived using radiative transfer theory by [ELIAS20011], and the result is also stated by [HHGTMS]. The notation used is shown in Figure 43.
@@
The ratio of the IOR of the coat dielectric ($n_{c}$) to that of the exterior medium ($n_{a}$) is denoted $\eta$. The Lambertian base albedo is denoted $\rho_{d}$. The dielectric layer absorption is parameterized by the transmittance along the normal direction, $T$. Light is incident from the upper hemisphere $\mathcal{H}_{+}$ at some incident angle cosine $\mu_{i}$. It enters the dielectric layer, bounces within, then exits back into $\mathcal{H}_{+}$ with some outgoing angle cosine $\mu_{o}$. Due to the azimuthal symmetry, the BRDF must be a function only of these two angle cosines, i.e., $f(\mu_{i},\mu_{o})$. By reciprocity we must have $f(\mu_{i},\mu_{o})=f(\mu_{o},\mu_{i})$. We assume that the incident light is unpolarized, and remains unpolarized after reflection and transmission (i.e., following the usual assumption in radiative transfer [^5] Though technically this is inconsistent, since even unpolarized incident light will become polarized after Fresnel transmission into the coat. This polarization is ignored and radiation assumed to be unpolarized, as in a standard RTE treatment of the problem., as employed by [ELIAS20011]). The formulas are all written assuming the spectral frequency is fixed.
We consider here only the case of a perfectly smooth dielectric coat. The extension to the rough case is briefy discussed by [HHGTMS].
In the presence of the Lambertian base, the full BRDF has the form [ELIAS20011]
@@
where $E_{F}(\eta)=2\int_{0}^{1}F_{R}(\mu,\eta)\,\mu\,\mathrm{d}\mu$ is the hemispherical (or average) albedo of the Fresnel factor.  [^6] A good approximation to $E_{F}(\eta)$ in the limited range $\eta\in[1,3]$ is given by [HHGTMS], who reports it to be accurate to within 0.15% over that IOR range: $E_{F}(\eta)\approx\ln\biggl(\frac{10893\eta-1438.2}{-774.4\eta^{2}+10212\eta+1}\biggr)\ .$ (124)
@@
The directional albedo thus simplifies to (as quoted by [HHGTMS]):
@@
Figure 47 compares the coat darkening factor $\mathbf{\Delta}$ in the cases of a rough and smooth base substrate (with a clear coat), for various coat IOR ratios $\eta$. A smooth base leads to less darkening (as well as less tinting in the absorptive case) since there is no TIR at all regardless of the angle of incidence, since all rays exit the coat at the same angle they entered (which couldn’t have been a TIR direction). In contrast, when diffuse scattering occurs, a very high proportion of rays end up undergoing TIR. This difference in the darkening of a coat for a smooth versus a rough base was also noted by [Weidlich2023].

### v2

@@ 5.3 Roughening
We intend to provide better approximations in the future. For example, a derivation of the roughening due to transmission through a rough GGX dielectric interface was provided by [7], which could provide a more physically accurate result.
@@ 5.4 Total internal reflection
We suggest as a reasonable practical approximation (following [37]) to simply invert the IOR ratio at the coat-base boundary, i.e., replace $\eta_{\mathrm{bc}}$ as follows:
@@
Both of these schemes produce a plausible appearance that eliminates the spurious TIR. A more physically correct approach would require something similar to the Weidlich–Wilkie layering model [53], or the layered material model described in [41], where the scattering through the layers is explicitly modeled via Monte Carlo simulation.
@@
Previous approaches include the microfacet-based sheen of [16] as used by the Autodesk Standard Surface model [17], and the single-scattering microflake model of Adobe Standard Material [37]. In OpenPBR, we adopt the multi-scattering volumetric microflake model of [56], which uses the SGGX microflake representation of [24] and provides efficient evaluation and importance sampling via linearly transformed cosines (LTCs) [25].
@@
  • Volumetric foundation: A homogeneous, purely scattering layer with a fiber-like SGGX microflake phase function [24], providing a physically plausible appearance.
@@
  • Layering model: The layer has unit optical thickness, is index-matched to the exterior medium (eliminating Fresnel at the interface), and conserves energy – all unreflected light transmits uncolored, as there is no absorption, which allows for principled layering (unlike the [16] model).
@@
where $\mathbf{F}$ = fuzz_color, $E_{\mathrm{fuzz}}(\mu_{o},\alpha)$ (termed $R$ by [56]) is the reflectance at angle cosine $\mu_{o}$ given roughness $\alpha$ = fuzz_roughness $\in[0,1]$, and $D(\mu_{i}|\mu_{o},\alpha)$ is a lobe defined by linear transformations of a cosine lobe (LTCs), where the transformation matrices (and $E_{\mathrm{fuzz}}$) are fitted (over $\mu_{o}$ and $\alpha$) to a simulation of the scattering in the volumetric fuzz microflake layer. Since the LTC lobe $D$ is a normalized PDF over the hemisphere, the resulting albedo of $f_{\mathrm{fuzz}}$ is $\mathbf{F}\,E_{\mathrm{fuzz}}(\mu_{o},\alpha)$.
@@
In the model of [56], the optical depth is at most 1, which means the opacity at normal incidence is at most $1-\exp(-1)=0.63$, i.e., at least 36% of the light gets through. In really thick dust, the opacity could be arbitrarily close to 1. If the model were to set the optical depth to a higher value (e.g., 3), thicker fuzz could be represented, though that would make the fuzz opacity more uniform across all viewing angles. Ideally, the optical depth would be exposed as a parameter. In other words, it would be helpful to be able to control the density and thickness in order to make a fully opaque layer of dust.
@@ 7.1 Thin-wall dielectric
In thin-walled mode, the translucent-base slab can be considered an infinitesimally thin sheet of dielectric (with an embedded absorbing but non-scattering medium), with the BSDF $f_{\mathrm{dielectric}}$ on both sides. A ladder of inter-reflections occurs inside this slab, producing a reflected lobe and undeflected refracted lobe. The transmission_color can be assumed to give the transmittance through the thin sheet at normal incidence due to absorption (with transmission_depth ignored). In the smooth case, the BRDF and BTDF of this sheet can be solved exactly by summing over a geometrical series of terms containing Fresnel and absorption factors, and this can be extended to a good approximation of the rough case by appropriately roughening the transmission lobe (as described by [35]). This model of thin-walled glass is a cheaper, much more convenient way to render windows than a finite thickness non-thin-walled mesh.
This model is similar to the ThinDielectricBxDF of PBRT, except with absorption and roughness accounted for [35, 41]. The roughening effect of the thin wall can be modeled by considering the broadening of the incident and transmitted rays. [7] provides a useful heuristic for this, but a more complete model would be a welcome advancement. Figure 39 shows the effect of varying the specular_roughness parameter on the appearance of a thin-wall glass material.
@@ 7.2 Thin-wall subsurface
It should also be noted that there currently isn’t a completely satisfactory model for the case of the described absorbing only thin-wall dielectric with rough boundaries, for example the ThinDielectricBxDF of PBRT [41] simplified to the case of smooth boundaries. There are suggested heuristics [35, 7] but a fully worked out model would be beneficial.
@@
  • There is interest in adding explicit support for a model for rendering glints [32, 55] or flakes [3], which are important effects for certain uses cases, such as car paint.
  • Development (currently in collaboration with [40]) of a set of example materials suitable as presets, and for demonstration of the model’s capabilities.
@@ Appendix B A mixture model example
In various practical models such as Disney’s “Principled” shader [10], Autodesk 3ds Max’s Physical Material [1] and Autodesk Standard Surface [17], the model is defined as a mixture (i.e., linear combination) of BSDF/BSSRDFs corresponding to the constituent BSDF lobes of the reflection and transmission. The light transport between the layers and overall energy balance is approximated via the mix weights. Given such a representation, the integration into a renderer is relatively straightforward as the individual lobes can be importance sampled according to their albedos and combined with direct lighting estimates via standard techniques such as multiple importance sampling (MIS) [51].
@@ Appendix C Similarity theory of subsurface anisotropy
As noted, there is the following “similarity relation” [58, 9] between the scattering coefficients in media with different anisotropies $g$ and $g^{\star}$:
@@ Appendix D F82-tint model fits for a variety of real metals
We present here fits to the F82-tint model used for the metallic slab. These were computed using publicly available tabulated IOR data [42]. Example programs for computing these fits can be found here:
@@ Appendix E Implementing thin‐film iridescence from first principles
Rather than the approach of [5] in which the thin-film interference effects were pre-integrated over color matching functions, we favour a more direct and general “locally spectral” approach that computes reflectance per light path by evaluating the full Fresnel and Airy interference stack – including complex amplitudes, polarizations, and phase shifts – at specific wavelengths sampled per path. This can begin with fixed red, green, and blue wavelengths, but better results are achieved by stochastically sampling wavelengths from approximate camera sensitivity curves (as described in Section 3.1.3). This enables convergence to neutral gray for very thick films and avoids the high-frequency color banding that fixed RGB wavelengths can produce.
The same wavelengths can also be reused to model dispersion, while all other BSDF components are free to ignore them and operate in RGB as usual. This approach uses only the Airy summation (Equation 3 from [5]) – a method whose origins trace back to 19th-century studies of thin-plate interference by G.B. Airy and others – but requires additional per-wavelength computations and assembling the necessary formulas from multiple sources rather than a single reference. We attempt to collect the necessary equations and implementation instructions here for convenience.
The geometry of the thin film is assumed to follow the model described in [5], as illustrated in Figure 41. Table 3 summarizes the relevant mathematical quantities. The film is taken to be applied at each microfacet of the base substrate, so can be assumed to be a flat slab with normal given by the local micronormal $\mathbf{m}$, and thickness $d$. The film is assumed to have real IOR $n_{2}$ and thickness $d$, and to be surrounded by ambient medium above with (real) IOR $n_{1}$, and substrate material with IOR $n_{3}$ (conducting or dielectric, where $n_{3}$ is complex in the conducting case). The incident light from direction $\boldsymbol{\omega}_{i}$ is at angle $\theta_{1}$ to $\mathbf{m}$, while the light refracted into the film makes angle $\theta_{2}$ with $\mathbf{m}$. Given this setup, we need to determine the albedo of the resulting configuration for incident light with wavelength $\lambda$, accounting for interference effects.
@@
Figure 41: Schematic of the thin-film model (following [5]).
@@ 3. Reciprocity for reverse direction
To avoid re-running the Fresnel equations, one can use the following symmetry relations between the reflection and transmission amplitudes [8]:
@@ Appendix F The physics of coat darkening
We present here some formulas for the physical darkening effect of light transport within a dielectric coat. In the case of a smooth dielectric coat on top of a Lambertian substrate (the so-called “interfaced Lambertian” model), there is an exact solution available, which we describe in detail below. This was derived using radiative transfer theory by [15], and the result is also stated by [13]. The notation used is shown in Figure 43.
@@
The ratio of the IOR of the coat dielectric ($n_{c}$) to that of the exterior medium ($n_{a}$) is denoted $\eta$. The Lambertian base albedo is denoted $\rho_{d}$. The dielectric layer absorption is parameterized by the transmittance along the normal direction, $T$. Light is incident from the upper hemisphere $\mathcal{H}_{+}$ at some incident angle cosine $\mu_{i}$. It enters the dielectric layer, bounces within, then exits back into $\mathcal{H}_{+}$ with some outgoing angle cosine $\mu_{o}$. Due to the azimuthal symmetry, the BRDF must be a function only of these two angle cosines, i.e., $f(\mu_{i},\mu_{o})$. By reciprocity we must have $f(\mu_{i},\mu_{o})=f(\mu_{o},\mu_{i})$. We assume that the incident light is unpolarized, and remains unpolarized after reflection and transmission (i.e., following the usual assumption in radiative transfer [^5] Though technically this is inconsistent, since even unpolarized incident light will become polarized after Fresnel transmission into the coat. This polarization is ignored and radiation assumed to be unpolarized, as in a standard RTE treatment of the problem., as employed by [15]). The formulas are all written assuming the spectral frequency is fixed.
We consider here only the case of a perfectly smooth dielectric coat. The extension to the rough case is briefy discussed by [13].
In the presence of the Lambertian base, the full BRDF has the form [15]
@@
where $E_{F}(\eta)=2\int_{0}^{1}F_{R}(\mu,\eta)\,\mu\,\mathrm{d}\mu$ is the hemispherical (or average) albedo of the Fresnel factor.  [^6] A good approximation to $E_{F}(\eta)$ in the limited range $\eta\in[1,3]$ is given by [13], who reports it to be accurate to within 0.15% over that IOR range: $E_{F}(\eta)\approx\ln\biggl(\frac{10893\eta-1438.2}{-774.4\eta^{2}+10212\eta+1}\biggr)\ .$ (124)
@@
The directional albedo thus simplifies to (as quoted by [13]):
@@
Figure 47 compares the coat darkening factor $\mathbf{\Delta}$ in the cases of a rough and smooth base substrate (with a clear coat), for various coat IOR ratios $\eta$. A smooth base leads to less darkening (as well as less tinting in the absorptive case) since there is no TIR at all regardless of the angle of incidence, since all rays exit the coat at the same angle they entered (which couldn’t have been a TIR direction). In contrast, when diffuse scattering occurs, a very high proportion of rays end up undergoing TIR. This difference in the darkening of a coat for a smooth versus a rough base was also noted by [54].
@@
  • v1.3 (Feb 14, 2026)
    • The coat darkening example code (Listing 2) was improved to account for the view-dependent absorption as well (and some typos in it were fixed).
    • Fixed broken alpha in some images.